OpenStudy (anonymous):

a baker can make h dozen of pastries in g hours. how man pastries can he make in b minutes?

OpenStudy (anonymous):

\[\left.\frac{12h}{g} \right/\text{hr} \]\[\left.\frac{12h}{60g} \right/\min \]